OpenStudy (anonymous):

Simplify the expression xy^2+3xy^2 6b^2c+(-b^2c) -y^3+7y^3 Please Help! Completely understood these the other day now i'm blank.

OpenStudy (zzr0ck3r):

xy^2+3xy^2 = 1xy^2+3xy^2 what is 3 of something plus 1 of something?

OpenStudy (zzr0ck3r):

1 apple + 3 apples = ? apples

OpenStudy (zzr0ck3r):

@BettyJean ?

OpenStudy (anonymous):

4 apso is it that simple though? that's whats really throwing me off My mind keeps making it more complicated like. so the first one would be 4xy^2 right?

OpenStudy (zzr0ck3r):

correct

OpenStudy (zzr0ck3r):

let a = xy^2 then you have a+3a = 4a

OpenStudy (zzr0ck3r):

=4xy^2

OpenStudy (zzr0ck3r):

6b^2c+(-b^2c) = 6b^2c-1b^2c = ?